. A marketing researcher conducted a survey of 25 shoppers randomly selected at the local mall and found that the average shopper intended to spend an average of $32.00 during their visit to the mall. The researcher also found that, on the average, the sample of shoppers surveyed spent $7.00 different than the mean of $32.00. (A) What is the value of the estimated standard error of the mean (SEM)?
Answer #1

Solution:

Standard error mean:
It is a standard deviation of the sampling distribution or estimate of standard deviation .
Mathematically, represented as σx = σ/√n
(A) We are given number of observations are surveyed = 25
and standard deviation = $7.00

σx = σ/√n = 7/√25
= 7/5
= 1.4